# Kondo breakdown in multi-orbital Anderson lattices induced by destructive hybridization interference

### v1

@@ Abstract
In this paper we consider a multi band extension to the periodic Anderson model. We use a single site DMFT(NRG) in order to study the impact of the conduction band mediated effective hopping of the correlated electrons between the correlated orbitals onto the heavy Fermi liquid formation. Whereas the hybridization of a single impurity model with two distinct conduction bands always adds up constructively, $T_{K}\propto\exp(-\alpha U/(\Gamma_{1}+\Gamma_{2}))$, we show that this does not have to be the case in lattice models, where, in remarkable contrast, also an emerging low-energy Fermi liquid scale $T_{0}\propto\exp(-\alpha U/(\Gamma_{1}-\Gamma_{2}))$ is possible due to quantum interference effects in multi band models. At high symmetry points, heavy Fermi liquid formation is suppressed which is associated with a breakdown of the Kondo effect. This results in an asymptotically scale-invariant (i.e., power-law) spectrum of the correlated orbitals $\propto|\omega|^{1/3}$ and a small Fermi surface including only the light quasi-particles. This orbital selective Mott phase demonstrates the possibility of metallic local criticality within the general framework of ordinary single site DMFT.
@@ 2.1 Model
$H^{\text{c}}$ $=\sum_{\mu}^{N_{\nu}}\sum_{\vec{k},\sigma}(\epsilon^{\text{c}}_{\vec{k}\nu}+\epsilon^{\text{c}}_{\nu})n^{\text{c}}_{k\sigma\nu}\,,$ (2)
@@
$H^{\text{hyb}}$ $=\sum_{\mu}^{N_{\nu}}\sum_{l,\sigma}V_{\nu}f^{\dagger}_{l\sigma}c_{l\sigma\nu}+\text{h.c.}\,,$ (4)
with $n^{\text{c}}_{\vec{k}\sigma\nu}=c^{\dagger}_{\vec{k}\sigma\nu}c_{\vec{k}\sigma\nu}$ and $n^{\text{f}}_{l\sigma}=f^{\dagger}_{l\sigma}f_{l\sigma}$. The creation and annihilation operators $c^{\dagger}_{\vec{k}\sigma\nu}$ and $c_{\vec{k}\sigma\nu}$ create (annihilate) electrons with spin $\sigma$ and momentum $\vec{k}$ in the non-interacting band with band index $\nu$ and dispersion $\epsilon^{\text{c}}_{\vec{k}\nu}+\epsilon^{\text{c}}_{\nu}$, while $f^{\dagger}_{l\sigma}$ ($f_{l\sigma}$) creates (annihilates) an electron with spin $\sigma$ in the $f$-orbital at the real space lattice site $l$. $N_{\nu}$ denotes the number of conduction bands and $N_{f}$ the number of correlated f-sites. The correlation effect of the $f$-orbitals is taken into account by the local Coulomb interaction of strength $U$. It is important to note that due to the hybridization $V_{\nu}$, the particle number in each band $\nu$ is not a conserved quantity, distinguishing it from multi-channel models with conserved channel symmetry [33, 34, 35].
@@ 2.2 Method
To tackle this complex model in the translational invariant form (PAM), we employ the DMFT [38, 39], which maps the lattice problem onto an effective single impurity Anderson model. The central DMFT equation therefore reads:
$G^{\text{f}}_{\text{lat},\sigma}(z)=G^{\text{f}}_{\text{imp},\sigma}(z)\,.$ (5)
Here, $G^{\text{f}}_{\text{lat},\sigma}(z)$ and $G^{\text{f}}_{\text{imp},\sigma}(z)$ describe the local single-particle $f$-Greens function of the lattice and the impurity problem, respectively. These spectral functions are formally defined as:
$G^{\text{f}}_{\text{lat},\sigma}(z)$ $=\sum_{\vec{k}}\frac{1}{z-\epsilon^{\text{f}}-\sum_{\nu}V_{\nu}^{2}G^{\text{0c}}_{\vec{k}\nu}(z)-\Sigma_{\sigma}(z)}\,,$ (6)
$G^{\text{f}}_{\text{imp},\sigma}(z)$ $=\frac{1}{z-\epsilon^{\text{f}}-\Delta_{\sigma}(z)-\Sigma_{\sigma}(z)}\,.$ (7)
In these equations, $G^{\text{0c}}_{\vec{k}\nu}(z)=(z-\epsilon^{\text{c}}_{\vec{k}\nu}-\epsilon^{\text{c}}_{\nu})^{-1}$ represents the propagator of a particle in the $\nu$-th non-interacting band, $\Sigma_{\sigma}(z)$ is the correlation-induced part of the self-energy, and $\Delta_{\sigma}(z)$ characterizes the effective medium in which the single impurity is embedded. Importantly, $\Sigma_{\sigma}(z)$ itself depends on the effective medium $\Delta_{\sigma}(z)$, making Eq. (5) a self-consistent problem. The only approximation made by DMFT is to neglect the momentum dependence of the self energy, $\Sigma_{\vec{k}\sigma}(z)\approx\Sigma_{\sigma}(z)$, which gets exact for a lattice with infinite coordination number, i.e. in infinite dimensions [38, 39].
@@ 3 Theoretical motivation
$\Delta_{lm}(\omega+i\delta)$ $=$ $\Re\Delta_{lm}(\omega)+i\Delta_{lm}(\omega)$ (9)
@@ 4 Results
To address the properties of the effective SIAM arising from the DMFT self-consistency equation, Eq. (5), we employ the Numerical Renormalization Group (NRG) technique, as implemented in the TRIQS open-source package [42]. Details on how we accurately solved the DMFT integrals can be found in appendix A.
@@ 4.1 Destructive hybridization interference
$\sum_{i}V_{\nu}^{2}G^{0\text{c}}_{\vec{k}\nu}(0)=\frac{1-\alpha}{1+\alpha}\frac{-V^{2}}{\epsilon^{\text{c}}_{\vec{k}1}+\epsilon^{\text{c}}_{1}}-i\pi V^{2}\delta(\epsilon^{\text{c}}_{\vec{k}1}+\epsilon^{\text{c}}_{1}),$ (16)
@@
Figure 1: Low energy scale $T_{0}$ as function of the dimensionless parameter $\gamma=\frac{1+\alpha}{1-\alpha}$ for different band widths $D$, $U/\Gamma=6$ and $\epsilon^{c}_{1}/D=0.6$. The dashed black line corresponds to the function $f(\gamma)=\exp(-\frac{\pi U}{8\Gamma}(\gamma-1))$.
Figure 1 displays the low energy scale $T_{0}(\gamma)/T_{0}(\gamma=1)\sqrt{\gamma}$ vs the dimensionless parameter $\gamma=\frac{1+\alpha}{1-\alpha}$ for various band widths, $U/\Gamma=6$ and $|\epsilon^{\text{c}}_{\nu}|/D=0.6$. The plots clearly demonstrate the exponential dependence of the low energy scale $T_{0}$ on $\gamma$. This corroborate our conjecture that the screening of the local moment in a periodic model is mainly driven by the real part of the conduction band propagator and is not dominated by the imaginary part. This surprising and somehow counterintuitive result is in contrast of the observation in the SIAM and the two-impurity Anderson model [46, 47].
By approximating the propagator at zero frequency ($\omega/D\approx 0$) in Eq. (16), we can determine the renormalization of the hybridization from its real part: $V^{2}(\gamma)=V^{2}/\gamma$. If we assume $\tilde{\Gamma}\approx\Gamma$ and substitute the renormalized hybridization into Eq. (13) for the single impurity Kondo temperature, the ratio of the low temperature scales should follow the relation
@@
Note that $\gamma=1$ implies $\alpha=0$ which corresponds to the conventional single band PAM. The factor $\gamma-1=2\alpha/(1-\alpha)$ is a measure how much the second band contributes to the local dynamics.
The r.h.s of Eq. (17) is added as a dashed black line to Figure 1. For small values of $\gamma-1$, we observe an excellent agreement between the ratio determined numerically with the DMFT(NRG) and analytic expression stated in Eq. (17). As $\gamma$ increases, the zero frequency estimate $V^{2}(\gamma)=V^{2}/\gamma$ tends to overestimate the renormalization. However, for larger band widths, the deviation from Eq. (17) becomes smaller. This behavior can be attributed to the fact that Eq. (14) is strictly valid at $\omega/D\approx 0$, and the frequency dependence of the propagator becomes more influential for small band widths $D$. For smaller band width, the finite frequencies start to contribute stronger which leads to deviations from the analytic estimate.
By tuning the parameter $\alpha$ from 0 to 1, we continuously increase the admixture of the second conduction band. $\alpha=0$ corresponds to the conventional single band PAM away from the Kondo insulator case, and $\alpha=1$ two a two band model with perfect destructive interference of the real part of the conduction band propagators. We demonstrated in Fig.  1 that the low temperature scale of the lattice, $T_{0}$, which has the meaning of an effective Kondo lattice scale, vanishes for $\alpha\to 1$ ($\gamma\to\infty$). The model undergoes a quantum phase transition: At the coupling $\alpha=1$, the Kondo screening breaks down and the effective impurity has a stable local moment (LM) FP.
@@ 4.3 Spectral functions and Kondo breakdown QCP
Having established the potential for destructive hybridization interference in the multi-orbital extension of the PAM, we will use our auxiliary two band model to focus on the special scenario characterized by symmetric couplings, $\alpha=V_{2}/V_{1}=1$, and $\gamma=\infty$, where Eq. (17) predicts a complete suppression of the low-energy screening scale. We systematically explored various parameter sets, encompassing diverse shapes for the $c$-orbital density of states (DOS), the ratio of $D_{1}/D_{2}$, and particle-hole asymmetric $f$-orbitals.
@@ 4.3.1 Particle-hole symmetric $f$-orbitals
Figure 2: Single-particle spectral function of (a) the $f$-orbitals, $\rho^{f}(\omega)$ (b) the $c_{1}$-orbitals, $\rho^{c}_{1}(\omega)$ and (c) the $c_{2}$-orbitals, $\rho^{c}_{2}(\omega)$, of the paramagnetic DMFT solution. The orange line represents the case where only the $c_{1}$ orbitals hybridize with the $f$-orbitals ($\alpha=0.0$), while the black line depicts results for identical hybridization with both bands ($\alpha=1.0$). The inset displays the $f$-spectra for small frequencies. The model parameters used are: $D/\Gamma=5$, $U/\Gamma=10$, and $|\epsilon^{\text{c}}_{\nu}|/D=0.4$.
@@
While the Hubbard bands in the $f$-spectra remain largely unchanged, the Kondo resonance has collapsed. This suppression of the Kondo resonance aligns with the prediction from Eq. (17), which anticipates the vanishing of $T_{0}$ when $\alpha=1$, effectively rendering $\gamma=\infty$. Indeed, as $\alpha$ increases from 0, the Kondo resonance rapidly (exponentially) narrows, exemplarily shown for $\alpha=0.5$ (orange lines).
@@ 4.3.2 Properties of the QCP for PH symmetric f-sites
Figure 3 illustrates the positive part of the different spectra on double logarithmic plots. Figure 3(a) displays the $c_{1}$-orbital spectra and Fig.  3(b) the $f$-orbital spectra. For better comparison, we plotted $\rho(\omega)/\rho(\omega_{0})$, with $\omega_{0}/\Gamma=10^{-5}$. As expected for the low-temperature physics of a QCP the spectra at high energies significantly differ at high energies for the different parameter sets but a universal low energy regime emerges. This low-frequency behavior adheres to a power law relationship, denoted as $\propto|\omega|^{r}$. We obtain the approximate exponents of $r^{\text{f}}\approx 0.33$ for the $f$-spectra and $r^{\text{c}}\approx-0.33$ for both $c$-spectra. Remarkably, this exponent remains unaffected by variations in the model parameters, emphasizing the concept of universality.
It’s important to note that the KB-QCP is independent of the Coulomb interaction strength, effectively rendering $U>U_{c}=0$. However, if we introduce a finite bandwidth also for the $f$-orbitals, $U_{c}$ is shifted to larger values, $U_{c}>0$. Within the OSM phase for $U>U_{c}$ we consistently observe the same universal behavior, consequently, our primary focus remains on the zero bandwidth limit.

### v4

@@ Abstract
In this paper we consider a multi band extension to the periodic Anderson model. We use a single site DMFT(NRG) in order to study the impact of the conduction band mediated effective hopping of the correlated electrons between the correlated orbitals onto the heavy Fermi liquid formation. Whereas the hybridization of a single impurity model with two distinct conduction bands always adds up constructively, $T_{K}\propto\exp(-\mathrm{const}\,U/(\Gamma_{1}+\Gamma_{2}))$, we show that this does not have to be the case in lattice models, where, in remarkable contrast, also an low-energy Fermi liquid scale $T_{0}\propto\exp(-\mathrm{const}\,U/(\Gamma_{1}-\Gamma_{2}))$ can emerge due to quantum interference effects in multi band models, where $U$ denotes the local Coulomb matrix element of the correlated orbitals and $\Gamma_{i}$ the local hybridization strength of band $i$. At high symmetry points, heavy Fermi liquid formation is suppressed which is associated with a breakdown of the Kondo effect. This results in an asymptotically scale-invariant (i.e., power-law) spectrum of the correlated orbitals $\propto|\omega|^{1/3}$, indicating non-Fermi liquid properties of the quantum critical point, and a small Fermi surface including only the light quasi-particles. This orbital selective Mott phase demonstrates the possibility of metallic local criticality within the general framework of ordinary single site DMFT.
@@ 2.1 Model
$H^{\text{c}}$ $=\sum_{\nu}^{N_{\nu}}\sum_{\vec{k},\sigma}(\epsilon^{\text{c}}_{\vec{k}\nu}+\epsilon^{\text{c}}_{\nu})n^{\text{c}}_{\vec{k}\sigma\nu}\,,$ (2)
@@
$H^{\text{hyb}}$ $=\sum_{\nu}^{N_{\nu}}\sum_{l,\sigma}V_{\nu}f^{\dagger}_{l\sigma}c_{l\sigma\nu}+\text{h.c.}\,,$ (4)
with $n^{\text{c}}_{\vec{k}\sigma\nu}=c^{\dagger}_{\vec{k}\sigma\nu}c_{\vec{k}\sigma\nu}$ and $n^{\text{f}}_{l\sigma}=f^{\dagger}_{l\sigma}f_{l\sigma}$. The creation and annihilation operators $c^{\dagger}_{\vec{k}\sigma\nu}$ and $c_{\vec{k}\sigma\nu}$ create (annihilate) electrons with spin $\sigma$ and momentum $\vec{k}$ in the non-interacting band with band index $\nu$ and dispersion $\epsilon^{\text{c}}_{\vec{k}\nu}+\epsilon^{\text{c}}_{\nu}$, while $c^{\dagger}_{l\sigma}$ ($c_{l\sigma}$) and $f^{\dagger}_{l\sigma}$ ($f_{l\sigma}$) create (annihilate) an electron with spin $\sigma$ in the $c$/$f$-orbital at the real space lattice site $l$. $N_{\nu}$ denotes the number of conduction bands and $N_{f}$ the number of correlated f-sites. The correlation effect of the $f$-orbitals is taken into account by the local Coulomb interaction of strength $U$. It is important to note that due to the hybridization $V_{\nu}$, the particle number in each band $\nu$ is not a conserved quantity, distinguishing it from multi-channel models with conserved channel symmetry [33, 34, 35].
@@ 2.2 Method
To tackle this complex model in the translational invariant form (PAM), we employ the DMFT [38, 39], which maps the lattice problem onto an effective single impurity Anderson model. The central DMFT equation, where we have dropped the spin index $\sigma$ for simplicity, reads:
$G^{\text{f}}_{\text{lat}}(z)\stackrel{{\scriptstyle!}}{{=}}G^{\text{f}}_{\text{imp}}(z)\,.$ (5)
Here, $G^{\text{f}}_{\text{lat}}(z)$ and $G^{\text{f}}_{\text{imp}}(z)$ describe the local single-particle $f$-Greens function of the lattice and the effective impurity problem, respectively. These spectral functions are formally defined as:
$G^{\text{f}}_{\text{lat}}(z)$ $=\sum_{\vec{k}}\left(z-\epsilon^{\text{f}}-\Delta^{\text{lat}}_{\vec{k}}(z)-\Sigma(z)\right)^{-1}\,,$ (6)
$G^{\text{f}}_{\text{imp}}(z)$ $=\left(z-\epsilon^{\text{f}}-\Delta^{\text{eff}}(z)-\Sigma(z)\right)^{-1}\,.$ (7)
In these equations $\Sigma(z)$ is the correlation-induced part of the self-energy, and $\Delta^{\text{eff}}(z)$ characterizes the effective medium in which the single impurity is embedded. The momentum dependent hybridization function $\Delta^{\text{lat}}_{\vec{k}}(z)$ is the physically relevant property, containing the information about the lattice model under investigation:
$\Delta^{\text{lat}}_{\vec{k}}(z)=\sum_{\nu}V_{\nu}^{2}G^{\text{0c}}_{\vec{k}\nu}(z),$ (8)
where $G^{\text{0c}}_{\vec{k}\nu}(z)=(z-\epsilon^{\text{c}}_{\vec{k}\nu}-\epsilon^{\text{c}}_{\nu})^{-1}$ represents the propagator of a particle in the $\nu$-th non-interacting band. From Eq. (5) one can formulate the DMFT self consistency equation as
$\Delta^{\text{eff}}(z)=z-\epsilon^{\text{f}}-\Sigma(z)-(G^{\text{f}}_{\text{lat}}(z))^{-1}\,.$ (9)
Importantly, $\Sigma(z)$ itself depends on the effective medium $\Delta^{\text{eff}}(z)$, making Eq. (9) a self-consistent problem. The only approximation made by DMFT is to neglect the momentum dependence of the self energy, $\Sigma_{\vec{k}}(z)\approx\Sigma(z)$, which gets exact for a lattice with infinite coordination number, i.e. in infinite dimensions [38, 39]. In order to get a self consistent solution we proceed as follows:
  1. (I)
    initialize the effective Medium: $\Delta^{\text{eff}}(z)=\sum_{\vec{k}}\Delta^{\text{lat}}_{\vec{k}}(z)$
  2. (II)
    solve the impurity problem defined by $\Delta^{\text{eff}}(z)$ to obtain the self-energy $\Sigma(z)$
  3. (III)
    calculate $G^{\text{f}}_{\text{lat}}(z)$ from Eq. (6) (see appendix A and Eq. (A.1) respectively)
  4. (IV)
    calculate $\Delta^{\text{eff}}(z)$ from Eq. (9)
  5. (V)
    continue steps (II) to (IV) until self-consistency is reached.
It is important to note, that while there always exists an equivalent single-band SIAM in case of a multi-band SIAM, this is generally not possible in the context of a multi-band extension to the PAM as discussed here. For more details on this topic, please refer to appendix B.
@@ 3 Theoretical motivation
$\Delta_{lm}(\omega+i\delta)$ $=$ $\Re\Delta_{lm}(\omega)+i\Im\Delta_{lm}(\omega)$ (11)
@@ 4 Results
To address the properties of the effective SIAM arising from the DMFT self-consistency equation, Eq. (5), we employ the Numerical Renormalization Group (NRG) technique, as implemented in the NRG-Lubljana interface [42, 43] to the TRIQS open-source package [44]. Details on how we accurately solved the DMFT integrals can be found in appendix A.
@@ 4.1 Destructive hybridization interference
$\sum_{\nu}V_{\nu}^{2}G^{0\text{c}}_{\vec{k}\nu}(0)=\frac{1-\alpha}{1+\alpha}\frac{-V^{2}}{\epsilon^{\text{c}}_{\vec{k}1}+\epsilon^{\text{c}}_{1}}-i\pi V^{2}\delta(\epsilon^{\text{c}}_{\vec{k}1}+\epsilon^{\text{c}}_{1}),$ (18)
@@
Figure 1: Low energy scale $T_{0}$ as function of the dimensionless parameter $\gamma(\alpha)=\frac{1+\alpha}{1-\alpha}$ for different band widths $D$, $U/\Gamma=6$ and $\epsilon^{c}_{1}/D=0.6$. The dashed black line corresponds to the function $f(\gamma)=\exp(-\frac{\pi U}{8\Gamma}(\gamma-1))$.
Figure 1 displays the low energy scale $T_{0}(\gamma)/T_{0}(\gamma=1)\sqrt{\gamma}$ vs the dimensionless parameter $\gamma(\alpha)=\frac{1+\alpha}{1-\alpha}$ for various band widths, $U/\Gamma=6$ and $|\epsilon^{\text{c}}_{\nu}|/D=0.6$. The plots clearly demonstrate the exponential dependence of the low energy scale $T_{0}$ on $\gamma$. This corroborate our conjecture that the screening of the local moment in a periodic model is mainly driven by the real part of the conduction band propagator and is not dominated by the imaginary part. This surprising and somehow counterintuitive result is in contrast of the observation in the SIAM and the two-impurity Anderson model [48, 49].
By approximating the propagator at zero frequency ($\omega/D\approx 0$) in Eq. (18), we can determine the renormalization of the hybridization from its real part: $V^{2}(\gamma)=V^{2}/\gamma$. If we substitute the renormalized hybridization into Eq. (15) for the single impurity Kondo temperature, the ratio of the low temperature scales should follow the relation
@@
Note that $\gamma(\alpha=0)=1$ corresponds to the conventional single band PAM, and $\gamma-1=2\alpha/(1-\alpha)$ is a measure how much the second band contributes to the local dynamics.
The r.h.s of Eq. (19) is added as a dashed black line to Figure 1. For small values of $\gamma-1$, we observe an qualitatively good agreement between the ratio determined numerically with the DMFT(NRG) and analytic expression stated in Eq. (19). As $\gamma$ increases, the zero frequency estimate $V^{2}(\gamma)=V^{2}/\gamma$ tends to overestimate the renormalization. However, for larger band widths, the deviation from Eq. (19) becomes smaller. This behavior can be attributed to the fact that Eq. (16) is strictly valid at $\omega/D\approx 0$, and the frequency dependence of the propagator becomes more influential for small band widths $D$. For smaller band width, the finite frequencies start to contribute stronger which leads to deviations from the analytic estimate.
By tuning the parameter $\alpha$ from 0 to 1, we continuously increase the admixture of the second conduction band. $\alpha=0$ corresponds to the conventional single band PAM away from the Kondo insulator case, and $\alpha=1$ two a two band model with perfect destructive interference of the real part of the conduction band propagators. We demonstrated in Fig.  1 that the low temperature scale of the lattice, $T_{0}$, which has the meaning of an effective Kondo lattice scale, vanishes for $\alpha\to 1$ ($\gamma(\alpha\to 1)\to\infty$). The model undergoes a quantum phase transition: At the coupling $\alpha=1$, the Kondo screening breaks down and the effective impurity has a stable local moment (LM) FP.
Please note that we leveraged a highly specific assumption regarding the second band, which mirrors the characteristics of the first band around the Fermi energy. This approach was chosen for its ability to provide a straightforward and lucid demonstration of the existence and significance of destructive hybridization interference. In essence, this phenomenon manifests whenever the absolute value of the sum of the real parts of individual contributions is smaller than the sum of the absolute values of these contributions:
$\left|\sum_{\nu}V_{\nu}^{2}\Re G^{0\text{c}}_{\vec{k}\nu}(0)\right|<\sum_{\nu}\left|V_{\nu}^{2}\Re G^{0\text{c}}_{\vec{k}\nu}(0)\right|.$ (20)
This condition doesn’t necessitate bands to be precisely inverse, rather, they should exhibit distinct curvatures around the Fermi energy. This broader context resembles our scenario where $V_{1}\neq V_{2}$ such that $\alpha\neq 1$.
@@ 4.3 Spectral functions and Kondo breakdown QCP
Having established the potential for destructive hybridization interference in the multi-orbital extension of the PAM, we will use our auxiliary two band model to focus on the special scenario characterized by symmetric couplings, $\alpha=V_{2}/V_{1}=1$, where Eq. (19) predicts a complete suppression of the low-energy screening scale. We systematically explored various parameter sets, encompassing diverse shapes for the $c$-orbital density of states (DOS), the ratio of $D_{1}/D_{2}$, and particle-hole asymmetric $f$-orbitals.
@@ 4.3.1 Particle-hole symmetric $f$-orbitals
Figure 2: Single-particle spectral function of (a) the $f$-orbitals, $\rho^{f}(\omega)$ (b) the $c_{1}$-orbitals, $\rho^{c}_{1}(\omega)$ and (c) the $c_{2}$-orbitals, $\rho^{c}_{2}(\omega)$, of the paramagnetic DMFT solution. The black line represents the case where only the $c_{1}$ orbitals hybridize with the $f$-orbitals ($\alpha=0.0$), while the red line depicts results for identical hybridization with both bands ($\alpha=1.0$). The inset displays the $f$-spectra for small frequencies. The model parameters used are: $D/\Gamma=5$, $U/\Gamma=10$, and $|\epsilon^{\text{c}}_{\nu}|/D=0.4$.
@@
While the Hubbard bands in the $f$-spectra remain largely unchanged, the Kondo resonance has collapsed. This suppression of the Kondo resonance aligns with the prediction from Eq. (19), which anticipates the vanishing of $T_{0}$ when $\gamma(\alpha=1)=\infty$. Indeed, as $\alpha$ increases from 0, the Kondo resonance rapidly (exponentially) narrows, exemplarily shown for $\alpha=0.5$ (orange lines).
@@ 4.3.2 Properties of the QCP for PH symmetric f-sites
Figure 3 illustrates the positive part of the different spectra on double logarithmic plots. Figure 3(a) displays the $c_{1}$-orbital spectra and Fig.  3(b) the $f$-orbital spectra. For better comparison, we plotted $\rho(\omega)/\rho(\omega_{0})$, with $\omega_{0}/\Gamma=10^{-5}$. As expected for the low-temperature physics of a QCP the spectra at high energies significantly differ for the different parameter sets but a universal low energy regime emerges. This low-frequency behavior adheres to a power law relationship, denoted as $\propto|\omega|^{r}$. We obtain the approximate exponents of $r^{\text{f}}\approx 0.33$ for the $f$-spectra and $r^{\text{c}}\approx-0.33$ for both $c$-spectra. Remarkably, this exponent remains unaffected by variations in the model parameters, emphasizing the concept of universality.
It’s crucial to emphasize that the KB-QCP remains unaffected by the strength of the Coulomb interaction, effectively rendering $U>U_{c}=0$. Indeed, this phase transition is discernible even in the band structure of the non-interacting model ($U=U_{c}=0$). Given three orbitals per unit cell, the band structure comprises three bands of specific widths that depend on model parameters. In the case of PH-symmetric $f$-orbitals, $\epsilon^{f}=U=0$, and $\alpha\to 1$, one of these bandwidths is gradually suppressed, resulting in a completely flat band precisely at the Fermi energy when $\alpha=1$. This observation isn’t unexpected, as the local moment FP of the SIAM is always continuously connected to the trivially decoupled FP for $V=0$. Thus, any transition from strong coupling to the local moment FP within single-site DMFT can be delineated by a non-interacting Gaussian FP.
When we introduce a finite bandwidth for the $f$-orbitals, the critical value $U_{c}$ shifts to larger magnitudes ($U_{c}>0$), eradicating the flat band in the non-interacting band structure. Nonetheless, the KB-QCP remains governed by a non-interacting FP. Throughout the OSM phase for $U>U_{c}$, we consistently observe the same universal behavior, consequently, our primary focus remains on the limit of zero $f$-bandwidth.